### v1

@@ 1 Introduction
Given the highly correlated physiological activities of the cardiovascular system, cardiovascular signals correspond to different modalities of a shared underlying process. Modeling their multi-modal conditional distributions can capture the relationships between available recorded signals (condition modalities) and desired generated signals (target modalities), thereby unifying potential generation tasks involving arbitrary signal modalities. In this work, we propose UniCardio, a specialized diffusion transformer designed to leverage multi-modal relationships among cardiovascular signals, ensuring effective and versatile signal generation (Fig. 1c). UniCardio adopts a transformer-based architecture with modality-specific encoders and decoders, alongside task-specific attention masks to regulate modality interactions. Different from other deep generative models goodfellow2020generative ; de2021diffusion , UniCardio captures multi-modal conditional distributions with a unified noise-to-data generative framework, mapping intra- and inter-modal relationships into a unified latent space for flexible use. To cope with varying combinations of condition and target modalities, UniCardio introduces a continual learning paradigm to incorporate generation tasks involving an increasing number of condition modalities, allowing the model to accommodate progressively complex relationships while explicitly balancing their contributions.
We pre-train UniCardio on the Cuffless BP dataset kachuee2016cuffless , which comprises 339 hours of trimodal recordings of PPG, ECG, and BP signals collected from ICU patients with diverse abnormal health conditions. Through joint utilization of multiple cardiovascular signals, UniCardio demonstrates state-of-the-art performance across a range of generation tasks, including denoising, imputation, and translation, outperforming recent task-specific baselines. The generated signals are implemented in a tuning-free manner to facilitate cardiovascular applications over multiple unseen domains, including detection of abnormal health conditions and estimation of vital signs, achieving comparable or even better performance than the ground-truth signals. The generated signals further ensure interpretability through displaying diagnostic characteristics of typical abnormalities, validated by clinician assessments. To our knowledge, UniCardio represents the first unified framework for cardiovascular signal generation, offering both technical innovations and practical insights to advance AI-assisted healthcare.
@@ 2.1 Unified multi-modal generative modeling for cardiovascular signals
To capitalize on the interrelated nature of cardiovascular signals for versatile generation, UniCardio fits their multi-modal conditional distributions to model the many-to-any relationships between the condition modalities and target modality, thereby unifying signal restoration and modality translation within a shared framework (Fig. 1b, Sec. 4.1). We obtain this objective with an advanced design of conditional diffusion models DDPM2020 ; SGMs2021 , which offer key advantages for generative modeling. These models operate through a forward process that gradually adds noise to the data, transforming it into a simple prior distribution (typically Gaussian), and a reverse process that learns to reconstruct the data by iteratively removing the noise. UniCardio develops an unconditional forward process that transforms different signal modalities into a unified prior, and a conditional reverse process that guides the iterative reconstruction of desired signals using conditional information alongside a unified diffusion step, allowing for flexible handling of diverse condition modalities in generation tasks (Fig. 1c, Sec. 4.2). The multi-modal conditional distributions are learned with transformer-based architectures bao2023one ; peebles2023scalable ; UViT2023 to accommodate various input-output configurations (Sec. 4.3, Supplementary Fig. 6), with key innovations including modality-specific encoders of multi-scale convolution, customized transformer modules with task-specific attention masks, and modality-specific decoders.
@@
Table 1: Comparison of UniCardio against task-specific baselines. We benchmark UniCardio on four well-studied yet challenging tasks, including PPG imputation, ECG imputation, PPG-to-ECG translation, and PPG-to-BP translation. We report the performance of UniCardio after the multi-modal pre-training, alongside two enhanced variants including UniCardio-F for further fine-tuning on the task of interest and UniCardio-M for incorporating more condition modalities during the testing stage. Performance metrics include RMSE, MAE, and KS-Test (lower values indicate better performance), averaged by 256 independent trials. The error bars represent the standard error of the mean. “-I” denotes the intermittent signals.
@@
Task           Method Input Signal RMSE ($\downarrow$) MAE ($\downarrow$) KS-Test ($\downarrow$) Model Size ($\downarrow$)
PPG Imputation PulseImpute xu2022pulseimpute PPG-I 0.0763 $\pm$ 0.0036 0.0571 $\pm$ 0.0023 0.1085 $\pm$ 0.0034 3.06M
DeepMVI bansal2021missing PPG-I 0.1745 $\pm$ 0.0063 0.1268 $\pm$ 0.0044 0.1395 $\pm$ 0.0042 2.78M
[…]
UniCardio-F PPG-I 0.0710 $\pm$ 0.0043 0.0515 $\pm$ 0.0028 0.0918 $\pm$ 0.0030 4.94M
UniCardio-M PPG-I, ECG, BP 0.0443 $\pm$ 0.0016 0.0347 $\pm$ 0.0012 0.0907 $\pm$ 0.0031 4.95M
ECG Imputation PulseImpute xu2022pulseimpute ECG-I 0.1391 $\pm$ 0.0033 0.1224 $\pm$ 0.0027 0.4679 $\pm$ 0.0089 3.06M
DeepMVI bansal2021missing ECG-I 0.2420 $\pm$ 0.0031 0.1395 $\pm$ 0.0022 0.2400 $\pm$ 0.0061 2.78M
[…]
UniCardio-F ECG-I 0.0938 $\pm$ 0.0045 0.0448 $\pm$ 0.0024 0.1199 $\pm$ 0.0037 4.94M
UniCardio-M PPG, ECG-I, BP 0.0385 $\pm$ 0.0027 0.0241 $\pm$ 0.0017 0.1024 $\pm$ 0.0032 4.95M
PPG-to-ECG Translation RDDM shome2024region PPG 0.5710 $\pm$ 0.0140 0.5155 $\pm$ 0.0153 0.7706 $\pm$ 0.0137 138.77M
CardioGAN sarkar2021cardiogan PPG 0.4313 $\pm$ 0.0100 0.3226 $\pm$ 0.0104 0.5208 $\pm$ 0.0146 5.97M
[…]
UniCardio-F PPG 0.1960 $\pm$ 0.0062 0.1165 $\pm$ 0.0059 0.2698 $\pm$ 0.0131 4.94M
UniCardio-M PPG, BP 0.1663 $\pm$ 0.0076 0.1199 $\pm$ 0.0070 0.3302 $\pm$ 0.0147 4.95M
PPG-to-BP Translation ABPNet paviglianiti2020noninvasive PPG 7.2994 $\pm$ 0.2725 5.6647 $\pm$ 0.2434 0.1745 $\pm$ 0.0071 1.59M
PPG2ABP ibtehaz2022ppg2abp PPG 6.1882 $\pm$ 0.2584 4.7835 $\pm$ 0.2076 0.1621 $\pm$ 0.0060 19.44M
[…]
UniCardio-F PPG 5.7924 $\pm$ 0.3250 4.5893 $\pm$ 0.2897 0.1567 $\pm$ 0.0079 4.94M
UniCardio-M PPG, ECG 6.5066 $\pm$ 0.3443 5.4908 $\pm$ 0.3077 0.2004 $\pm$ 0.0086 4.95M
@@ 2.2 Versatile high-quality cardiovascular signal generation
As shown in Table 1, UniCardio after multi-modal pre-training has already achieved competing or even better performance across all tasks compared to the corresponding baselines. This explicitly demonstrates the benefits of leveraging multi-modal relationships during the pre-training stage. We further evaluate two variants of UniCardio to explore additional enhancements: The first variant UniCardio-F fine-tunes the model using previous pre-training data for a specific generation task, making it more tailored to the task of interest. The second variant UniCardio-M employs more available condition modalities during the testing stage to leverage their complementary information. Both variants deliver substantial improvements, outperforming all baselines by a large margin. Moreover, UniCardio performs all tasks with a comparably small amount of parameters, where the use of different modalities only requires adding the corresponding encoders and decoders (around 0.004M for each modality), making it applicable to wearable devices. In contrast, implementing these functions with task-specific methods would require combining multiple proprietary models, leading to tens of times greater parameter overhead. Such efficiency underscores the particular benefits of incorporating multi-modal relationships within a unified framework, as UniCardio does.
@@ 2.3 Robust real-time health monitoring with UniCardio
UniCardio’s integrated functions of signal restoration and modality translation provide a comprehensive enhancement to cardiovascular signal processing (Fig. 1). To demonstrate its practical effectiveness, we apply UniCardio to the publicly available datasets of unseen domains and explore representative applications spanning two major areas: detecting abnormal health conditions and estimating vital signs. Depending on the specific characteristics of each dataset (e.g., availability of multi-modal signals or disease annotations) as well as the unique demands of each application, we assess UniCardio’s generative capabilities in a targeted manner, centering around ECG as a typical example in cardiovascular diagnostics. All evaluations are conducted in a tuning-free manner, enabling the pre-trained generative model to be directly applied to diverse healthcare scenarios of unseen domains without additional fine-tuning.
Figure 4: Versatile generation assisted cardiovascular applications. a-c, Detection of ST change using the PTBXL dataset wagner2020ptb . d-f, Detection of hypertrophy using the PTBXL dataset wagner2020ptb . g-i, Detection of atrial fibrillation (AF) using the MIMIC PERform AF dataset charlton2022detecting . j, Estimation of heart rate (HR) using the WESAD dataset schmidt2018introducing . k-l, Estimation of diastolic BP and systolic BP using unseen data of the Cuffless BP dataset kachuee2016cuffless . “-N” indicates noisy signals. “-I” indicates intermittent signals. “-G” indicates generated signals. The generated signals are produced by UniCardio in a tuning-free manner. Here we do not include the error bar since most cases are averaged over the entire test set without repeated sampling.
We first evaluate the detection of multiple cardiovascular abnormalities with the PTBXL dataset wagner2020ptb , which includes only ECG signals annotated with abnormal health conditions. Of these, we adopt the ST change (Fig. 4a-c) and hypertrophy (Fig. 4d-f) as representative examples. This scenario focuses on UniCardio’s denoising capability, which mitigates noise-induced degradation during recording and transmission. The noisy signals exhibit low accuracy, extremely low specificity, and inflated sensitivity due to excessive false-positive classifications. In contrast, the denoised signals generated by UniCardio resolve these issues, achieving performance comparable to the ground-truth signals. Additionally, missing segments that may result from sensor disconnections can moderately affect both abnormalities, which are addressed by UniCardio’s imputation capability.
@@
Next, we evaluate the estimation of vital signs, specifically heart rate (HR) using the WESAD dataset schmidt2018introducing (Fig. 4j), and systolic and diastolic BP using unseen data from the Cuffless BP dataset kachuee2016cuffless (Fig. 4k-l). Both datasets encompass multi-modal cardiovascular signals, allowing to validate the complementary benefits of generating non-wearable signals from wearable devices. For HR estimation, UniCardio can effectively translate high-quality ECG signals from wearable PPG signals and impute intermittent ECG signals to restore missing segments. The generated ECG signals achieve significantly more accurate HR estimation compared to the original PPG signals and intermittent ECG signals, substantially reducing the mean absolute error (MAE). For BP estimation, we focus on PPG-to-ECG translation, as intermittent signals have limited impact on systolic and diastolic BP. Combining the original PPG signals with the generated ECG signals markedly improves BP estimation, with a notable reduction in MAE for both systolic and diastolic values.
@@
To address the efficiency concerns commonly raised with diffusion models, we validate that UniCardio maintains comparable performance even with significantly reduced sampling steps (Supplementary Sec. D and Table 3). Such efficient sampling creates an average delay of only 0.4 seconds, making UniCardio well-suited for real-time monitoring. Together, these advantages establish UniCardio as a practical, robust, and interpretable tool for advancing cardiovascular healthcare.

### v2

@@ 1 Introduction
Given the highly correlated physiological activities of the cardiovascular system, cardiovascular signals correspond to different modalities of a shared underlying process. Modeling their multi-modal conditional distributions can capture the relationships between available recorded signals (condition modalities) and desired generated signals (target modalities), thereby unifying potential generation tasks involving arbitrary signal modalities. In this work, we propose UniCardio, a specialized diffusion transformer designed to leverage multi-modal relationships among cardiovascular signals, ensuring effective and versatile signal generation (Fig. 1c). UniCardio adopts a transformer-based architecture with modality-specific encoders and decoders, alongside task-specific attention masks to regulate modality interactions. Different from other deep generative models goodfellow2020generative ; de2021diffusion , UniCardio captures multi-modal conditional distributions with a unified noise-to-data generative framework, mapping intra- and inter-modal relationships into a unified latent space for flexible use. To cope with varying combinations of condition and target modalities, UniCardio introduces a continual learning paradigm to incorporate generation tasks involving an increasing number of condition modalities, thereby accommodating progressively complex relationships while explicitly balancing their contributions.
We pre-train UniCardio on the Cuffless BP dataset kachuee2016cuffless , which comprises 339 hours of trimodal recordings of PPG, ECG, and BP signals collected from ICU patients with diverse abnormal health conditions. Through joint utilization of multiple cardiovascular signals, UniCardio demonstrates state-of-the-art performance across a range of generation tasks, including denoising, imputation, and translation, outperforming recent task-specific baselines. The generated signals have promoted downstream cardiovascular applications over multiple unseen domains, including detection of abnormal health conditions and estimation of vital signs, achieving comparable or even better performance than the ground-truth signals. The generated signals further ensure interpretability through displaying diagnostic characteristics of typical abnormalities, validated by clinician assessments. To our knowledge, UniCardio represents the first unified framework for cardiovascular signal generation. It integrates technical innovations in multi-modal generation and continual learning into practical solutions for signal restoration and modality translation, serving as a versatile foundation for advancing real-time health monitoring.
@@ 2.1 Unified multi-modal generative modeling for cardiovascular signals
To capitalize on the interrelated nature of cardiovascular signals for versatile generation, UniCardio fits their multi-modal conditional distributions to model many-to-any relationships between condition modalities and target modality, thereby unifying signal restoration and modality translation within a shared framework (Fig. 1b, Sec. 4.1). We obtain this objective with an advanced design of conditional diffusion models DDPM2020 ; SGMs2021 , which offer key advantages for generative modeling. These models operate through a forward process that gradually adds noise to the data, transforming it into a simple prior distribution (typically Gaussian), and a reverse process that learns to reconstruct the data by iteratively removing the noise. UniCardio develops an unconditional forward process that transforms different signal modalities into a unified prior, and a conditional reverse process that guides the iterative reconstruction of desired signals using conditional information alongside a unified diffusion step, enabling flexible handling of diverse condition modalities in generation tasks (Fig. 1c, Sec. 4.2). The conditional distributions are learned with transformer-based architectures bao2023one ; peebles2023scalable ; UViT2023 to accommodate various input-output configurations (Sec. 4.3, Supplementary Fig. 1), with key innovations including modality-specific encoders of multi-scale convolution, customized transformer modules with task-specific attention masks, and modality-specific decoders.
@@
Table 1: Comparison of UniCardio against task-specific baselines. We benchmark UniCardio on four well-studied yet challenging tasks, including PPG imputation, ECG imputation, PPG-to-ECG translation, and PPG-to-BP translation. We report the performance of UniCardio after the multi-modal pre-training, alongside two enhanced variants including UniCardio-F for further fine-tuning on the task of interest and UniCardio-M for incorporating more condition modalities during the testing stage. Performance metrics include RMSE, Min-RMSE, MAE, Min-MAE, and KS-Test (lower values indicate better performance), averaged by 256 independent trials. The error bars represent the standard error of the mean. “-I” denotes the intermittent signals. The best performance for each task is highlighted in bold.
@@
Task           Method Input Signal RMSE ($\downarrow$) Min-RMSE ($\downarrow$) MAE ($\downarrow$) Min-MAE ($\downarrow$) KS-Test ($\downarrow$) Model Size ($\downarrow$)
PPG
Imputation
PulseImpute xu2022pulseimpute PPG-I 0.0763 $\pm$ 0.0036 0.0393 $\pm$ 0.0018 0.0571 $\pm$ 0.0023 0.0164 $\pm$ 0.0007 0.1085 $\pm$ 0.0034 3.06M
DeepMVI bansal2021missing PPG-I 0.1745 $\pm$ 0.0063 0.0890 $\pm$ 0.0030 0.1268 $\pm$ 0.0044 0.0382 $\pm$ 0.0013 0.1395 $\pm$ 0.0042 2.78M
[2pt/2pt] UniCardio PPG-I 0.1146 $\pm$ 0.0063 0.0585 $\pm$ 0.0031 0.0797 $\pm$ 0.0043 0.0244 $\pm$ 0.0013 0.1084 $\pm$ 0.0039 5.51M
UniCardio-F PPG-I 0.0710 $\pm$ 0.0043 0.0226 $\pm$ 0.0005 0.0515 $\pm$ 0.0028 0.0104 $\pm$ 0.0002 0.0918 $\pm$ 0.0030 5.51M
UniCardio-M PPG-I, ECG, BP 0.0443 $\pm$ 0.0016 0.0243 $\pm$ 0.0009 0.0347 $\pm$ 0.0012 0.0104 $\pm$ 0.0004 0.0907 $\pm$ 0.0031 6.09M
ECG
Imputation
PulseImpute xu2022pulseimpute ECG-I 0.1391 $\pm$ 0.0033 0.0760 $\pm$ 0.0018 0.1224 $\pm$ 0.0027 0.0367 $\pm$ 0.0008 0.4679 $\pm$ 0.0089 3.06M
DeepMVI bansal2021missing ECG-I 0.2420 $\pm$ 0.0031 0.1267 $\pm$ 0.0017 0.1395 $\pm$ 0.0022 0.0413 $\pm$ 0.0007 0.2400 $\pm$ 0.0061 2.78M
[2pt/2pt] UniCardio ECG-I 0.1756 $\pm$ 0.0042 0.0891 $\pm$ 0.0025 0.0750 $\pm$ 0.0026 0.0251 $\pm$ 0.0009 0.1486 $\pm$ 0.0055 5.51M
UniCardio-F ECG-I 0.0938 $\pm$ 0.0045 0.0622 $\pm$ 0.0022 0.0448 $\pm$ 0.0024 0.0177 $\pm$ 0.0006 0.1199 $\pm$ 0.0037 5.51M
UniCardio-M PPG, ECG-I, BP 0.0385 $\pm$ 0.0027 0.0210 $\pm$ 0.0015 0.0241 $\pm$ 0.0017 0.0073 $\pm$ 0.0005 0.1024 $\pm$ 0.0032 6.09M
PPG-to-ECG
Translation
RDDM shome2024region PPG 0.5710 $\pm$ 0.0140 0.5233 $\pm$ 0.0148 0.5155 $\pm$ 0.0153 0.4839 $\pm$ 0.0153 0.7706 $\pm$ 0.0137 138.77M
CardioGAN sarkar2021cardiogan PPG 0.4313 $\pm$ 0.0100 0.3105 $\pm$ 0.0103 0.3226 $\pm$ 0.0104 0.2598 $\pm$ 0.0103 0.5208 $\pm$ 0.0146 5.97M
[2pt/2pt] UniCardio PPG 0.2747 $\pm$ 0.0067 0.2173 $\pm$ 0.0069 0.1937 $\pm$ 0.0070 0.1668 $\pm$ 0.0070 0.4407 $\pm$ 0.0154 5.51M
UniCardio-F PPG 0.1960 $\pm$ 0.0062 0.1326 $\pm$ 0.0056 0.1165 $\pm$ 0.0059 0.0888 $\pm$ 0.0056 0.2698 $\pm$ 0.0131 5.51M
UniCardio-M PPG, BP 0.1663 $\pm$ 0.0076 0.1396 $\pm$ 0.0071 0.1199 $\pm$ 0.0070 0.1091 $\pm$ 0.0068 0.3302 $\pm$ 0.0147 6.09M
PPG-to-BP
Translation
ABPNet paviglianiti2020noninvasive PPG 7.2994 $\pm$ 0.2725 6.1959 $\pm$ 0.2566 5.6647 $\pm$ 0.2434 4.8927 $\pm$ 0.2336 0.1745 $\pm$ 0.0071 1.59M
PPG2ABP ibtehaz2022ppg2abp PPG 6.1882 $\pm$ 0.2584 4.9102 $\pm$ 0.2419 4.7835 $\pm$ 0.2076 3.4902 $\pm$ 0.2000 0.1621 $\pm$ 0.0060 19.44M
[2pt/2pt] UniCardio PPG 10.1538 $\pm$ 0.4213 8.8916 $\pm$ 0.3937 8.3721 $\pm$ 0.3889 7.5386 $\pm$ 0.3690 0.2668 $\pm$ 0.0097 5.51M
UniCardio-F PPG 5.7924 $\pm$ 0.3250 4.5986 $\pm$ 0.2993 4.5893 $\pm$ 0.2897 3.8438 $\pm$ 0.2755 0.1567 $\pm$ 0.0079 5.51M
UniCardio-M PPG, ECG 6.5066 $\pm$ 0.3443 5.9460 $\pm$ 0.3206 5.4908 $\pm$ 0.3077 5.0413 $\pm$ 0.2923 0.2004 $\pm$ 0.0086 6.09M
@@ 2.2 Versatile high-quality cardiovascular signal generation
To further investigate the benefit of joint modeling, we conduct an ablation study: Baseline-1, single-modality models trained on only one-condition imputation tasks; Baseline-2, dual-modality models trained on one-condition translation tasks and two-condition imputation tasks; and Baseline-3, tri-modality models trained on two-condition translation tasks and three-condition imputation tasks. All models were trained for the same number of effective epochs to ensure fairness. As shown in Supplementary Table 2, UniCardio achieves comparable or slightly inferior performance to Baseline-1 on one-condition imputation and Baseline-2 on one-condition translation, despite being a compact model that handles a much broader set of tasks across all modality combinations. Notably, UniCardio significantly outperforms Baseline-2 on two-condition imputation tasks (31.69% RMSE ratio), and outperforms Baseline-3 on two-condition translation tasks (35.97% RMSE ratio) and three-condition imputation tasks (63.85% RMSE ratio). These results highlight knowledge transfer across task-modality combinations: training on tasks involving fewer modalities facilitates tasks involving more modalities.
@@
As shown in Table 1, UniCardio after multi-modal pre-training has already achieved competing or even better performance across all tasks compared to the corresponding baselines. We consider UniCardio’s two variants to explore additional enhancements: The first variant UniCardio-F fine-tunes the model using previous pre-training data for a specific generation task, making it more tailored to the task of interest. The second variant UniCardio-M employs more available condition modalities during the testing stage to leverage their complementary information. Both variants deliver substantial improvements, largely outperforming all baselines in most cases. To provide additional reference, we also evaluate classical signal processing baselines, such as a wavelet denoising algorithm sendur2003bivariate for denoising and a Gerchberg–Papoulis algorithm gerchberg1974super ; papoulis2003new for imputation (Supplementary Table 3). As expected, these traditional methods clearly underperform UniCardio as well as other deep learning–based methods.
Moreover, UniCardio performs all tasks with a comparably small amount of parameters, where the use of different modalities only requires adding the corresponding encoders and decoders (around 0.3M for each modality), making it applicable to wearable devices. In contrast, implementing these functions with task-specific methods would require combining multiple proprietary models, leading to tens of times greater parameter overhead. Such efficiency underscores the particular benefits of incorporating multi-modal relationships within a unified framework, as UniCardio does.
@@ 2.3 Robust real-time health monitoring with UniCardio
UniCardio’s integrated functions of signal restoration and modality translation provide a comprehensive enhancement to cardiovascular signal processing (Fig. 1). To demonstrate its practical effectiveness, we apply UniCardio to the publicly available datasets of unseen domains and explore representative applications spanning two major areas: detecting abnormal health conditions and estimating vital signs. Unless otherwise specified, we implement the generated signals without additional fine-tuning. Depending on the specific characteristics of each dataset (e.g., availability of multi-modal signals or disease annotations) as well as the unique demands of each application, we assess UniCardio’s generative capabilities in a targeted manner, centering around ECG as a typical example in cardiovascular diagnostics.
Figure 4: Versatile generation assisted cardiovascular applications. a-c, Detection of ST change using the PTBXL dataset wagner2020ptb . d-f, Detection of hypertrophy using the PTBXL dataset wagner2020ptb . g-i, Detection of atrial fibrillation (AF) using the MIMIC PERform AF dataset charlton2022detecting . j, Estimation of heart rate (HR) using the WESAD dataset schmidt2018introducing . k-l, Estimation of systolic BP (SBP) and diastolic BP (SBP) using the MIMIC dataset moody1996database . “-N” indicates noisy signals. “-I” indicates intermittent signals. “-G” indicates generated signals. “-Ave” indicates using the average SBP/DBP of the training set as the prediction. Here we do not include the error bar since most cases are averaged over the entire test set without repeated sampling.
We first evaluate the detection of multiple cardiovascular abnormalities using the PTBXL dataset wagner2020ptb , which includes only ECG signals annotated with abnormal health conditions. Of these, we adopt the ST change (Fig. 4a-c) and hypertrophy (Fig. 4d-f) as representative examples. This scenario focuses on UniCardio’s denoising capability, which mitigates noise-induced degradation during recording and transmission. The noisy signals exhibit low accuracy, extremely low specificity, and inflated sensitivity due to excessive false-positive classifications. In contrast, the denoised signals generated by UniCardio resolve these issues, achieving performance comparable to the ground-truth signals. Additionally, missing segments that may result from sensor disconnections can moderately affect both abnormalities, which are addressed by UniCardio’s imputation capability.
@@
Next, we evaluate the estimation of vital signs, focusing on heart rate (HR) using the WESAD dataset schmidt2018introducing and systolic/diastolic BP (SBP/DBP) using the MIMIC dataset moody1996database . These multi-modal datasets allow us to examine the benefits of obtaining non-wearable signals from wearable devices. For HR estimation, UniCardio can effectively translate high-quality ECG signals from wearable PPG signals and impute intermittent ECG signals to restore missing segments, substantially reducing the mean absolute error (MAE) compared to the original PPG and degraded ECG signals, respectively (Fig. 4j). For BP estimation, we focus on translation tasks, as intermittent signals have limited impact on SBP/DBP prediction. Although UniCardio pretrained on Cuffless BP can be directly applited to MIMIC and generate high-quality waveforms (Supplementary Fig. 2), we find that fine-tuning is necessary to obtain accurate SBP/DBP values. The fine-tuned UniCardio achieves consistent improvements over baseline models that predict SBP/DBP directly from PPG and ECG on a strictly held-out MIMIC test set with no subject overlap (Fig. 4k,l). We also consider a naive baseline that simply uses the average SBP/DBP from the training set for prediction, which likewise underperforms ours (Fig. 4k,l).
Interestingly, we observe that UniCardio’s ECG-to-BP translation consistently outperforms PPG-to-BP translation on both Cuffless BP (Fig. 3g) and MIMIC (Supplementary Fig. 2 and Fig. 4k,l). Although prior studies have reported that BP estimation from ECG is generally less accurate than from PPG finnegan2023features ; landry2023current , we hypothesize that UniCardio’s multi-task, multi-modal pre-training facilitates shared representation learning across diverse task-modality combinations, enabling it to extract latent predictive features from ECG that traditional models underutilize. We defer a systematic examination of this phenomenon across broader patient cohorts and clinical contexts to future studies.
@@
To address the efficiency concerns commonly raised with diffusion models, we validate that UniCardio maintains comparable performance even with significantly reduced sampling steps (Supplementary Sec. D and Table 4). This efficient sampling results in an average inference delay of less than only 0.4 seconds per 4-second segment, well within the constraints of real-time monitoring. The accompanying pre-processing pipeline (including signal normalization, segment extraction, and entropy-based filtering) adds only 0.0034 seconds per segment on a standard CPU of the same server, rendering its overhead negligible. Together, these advantages establish UniCardio as a practical, robust, and interpretable tool for advancing cardiovascular healthcare.